# Stabilizing Policy Gradient Methods via Reward Profiling

### v1

@@ 6.2 Experiment with Multi-robot Task
For a realistic and domain-specific evaluation, we deployed it in a multi-agent robotic setting on the Unity ML-Agents Reacher task (juliani2020; cohen2022). Unity provides a flexible, visually realistic, and customizable simulation environment; see Figure 3. Here, the environment simulates 20 robotic arms operating in parallel, each receiving a 33-dimensional state vector (joint angles, velocities, and target position) and outputting 4D continuous torques. Agents earn a dense reward of +0.1+0.1 per timestep for maintaining their fingertip in a moving target zone. We plugged our profiling wrapper (the TP variant) into the platform atop a decentralized DDPG architecture, with policy updates evaluated every 50 episodes through 5-agent consensus voting.
Figure 5: Sensitivity to evaluation rollouts $E$ for DDPG+Three-Points on HopperBulletEnv-v0. Curves show mean selected returns (±1 std). Small EE(10–20) yields noisy, unstable updates; large EE(200) improves stability but slows progress; mid-range EE(50–100) balances both.
@@ 7 Evaluation Rollout Sensitivity
Using an NVIDIA V100 GPU, baseline training of 100K steps took 10–30 min, with the TP variant adding only about 20% wall-clock time. Figure 5 shows how the evaluation budget shapes learning. At small $E$ (10-20), the Monte Carlo estimates exhibit noisy, spurious updates and wide confidence bands. This results in large fluctuations and wide confidence bands. At the other extreme, very large $E$ (e.g., 200), the algorithm advances only slowly once the noise floor is suppressed. Intermediate values (E=50E=50–100) suppress the worst of the noise without over‑constraining the update rule, yielding both rapid early gains and a much narrower variance envelope. So roughly, pushing $E$ well beyond this critical scale offers diminishing returns. Once the estimation error is negligible, further rollouts merely add computational cost and slow down true policy updates. This reflects the batch-size trade‑off in SGD (mccandlish2018empirical), where a task‑dependent critical batch size marks the point beyond which additional samples no longer improve data efficiency. By analogy, one can think of a critical evaluation budget $E_{\rm crit}$, and choose $E\approx E_{\rm crit}$, to optimally balance variance reduction against responsiveness. Our framework complements other recent monotonic improvement goals (xie2025spo) without relying on stochastic density ratios, hence applies uniformly to both stochastic and deterministic settings.
@@ A.1   Concentration Inequality (Proof of Lemma 1)
Applying Hoeffding’s inequality to the bounded, independent variables $\{X_{i}\}$ (hoeffding1994probability) gives for any $\epsilon>0$ the one‑sided tail bounds
@@ A.4   Smoothness and Descent Properties
Follows from standard policy gradient smoothness results (e.g. Lemma D.3 in (agarwal2021theory)), using the bounded Fisher information in Assumption 1. ∎
@@ A.5   Proof of Theorem 1
so $\min_{t<T}\|\nabla J(\pi_{\theta_{t}})\|=O(T^{-1/4})$. Lemma 41̇ of(agarwal2021theory) then gives
@@ A.6   Proof of Theorem 2 (Biased Critic Case)
where $\epsilon_{t}$ is the bias that we assume is also upper bounded as well as its gradient by some $\epsilon\geq 0$, i.e., the estimation of the value function is accurate in both zeroth and first order. Such an assumption can be satisfied when a deep neural network is used to estimate value functions, e.g., (du2019gradient; neyshabur2017implicit; miyato2018spectral), and is widely assumed in actor-critic analysis, e.g., (wang2019neural; zhou2023single; chen2023global; zhang2020provably; qiu2021finite; kumar2023sample; xu2020non; xu2020improving; suttle2023beyond).
@@
(Lemma D.3 of (agarwal2021theory)) The value functions are $L$-smooth, i.e., there exists some constant $L$, such that
@@
Following a similar vanilla SGD analysis (dutta2020discrepancy; dutta2023demystifying), we obtain the results.
@@ B.4   Additional Experiments
In the main paper (Figure 2), we showed four continuous‐control suites for PPO, TRPO, and DDPG. Here, we complete the picture by adding results on two Box2D tasks (CarRacing, BipedalWalker) and two high‐DOF Bullet tasks (HalfCheetah, Humanoid). We also evaluated our profiling wrapper on the state-of-the-art off-policy actor-critic, TD3 (fujimoto18addressing), that extends DDPG with clipped-double Q-learning, delayed policy updates, and target-policy smoothing. All of TD3’s core improvements—reduced overestimation bias from twin critics, more accurate actor gradients via delayed updates, and smoother targets via policy noise—are preserved inside our wrapper while reward profiling yields similar stability gains across both on‐ and off‐policy methods.

### v2

@@ 6.2 Experiment with Multi-robot Task
For a realistic and domain-specific evaluation, we deployed it in a multi-agent robotic setting on the Unity ML-Agents Reacher task (Juliani et al. 2020; Cohen et al. 2022). Unity provides a flexible, visually realistic, and customizable simulation environment; see Figure 3. Here, the environment simulates 20 robotic arms operating in parallel, each receiving a 33-dimensional state vector (joint angles, velocities, and target position) and outputting 4D continuous torques. Agents earn a dense reward of +0.1+0.1 per timestep for maintaining their fingertip in a moving target zone. We plugged our profiling wrapper (the TP variant) into the platform atop a decentralized DDPG architecture, with policy updates evaluated every 50 episodes through 5-agent consensus voting.
Figure 5: Sensitivity to evals $E$ for DDPG+TP on Hopper. Smaller EE(10–20) causes noisy, unstable updates; large EE(200) improves stability but slows progress; mid-range EE(50–100) balances both.
@@ 7 Evaluation Rollout Sensitivity
Using an NVIDIA V100 GPU, baseline training of 100K steps took 10–30 min, with the TP variant adding only about 20% wall-clock time. Figure 5 shows how the evaluation budget shapes learning. At small $E$ (10-20), the Monte Carlo estimates exhibit noisy, spurious updates and wide confidence bands. This results in large fluctuations and wide confidence bands. At the other extreme, very large $E$ (e.g., 200), the algorithm advances only slowly once the noise floor is suppressed. Intermediate values (E=50E=50–100) suppress the worst of the noise without over‑constraining the update rule, yielding both rapid early gains and a much narrower variance envelope. So roughly, pushing $E$ well beyond this critical scale offers diminishing returns. Once the estimation error is negligible, further rollouts merely add computational cost and slow down true policy updates. This reflects the batch-size trade‑off in SGD (McCandlish et al. 2018), where a task‑dependent critical batch size marks the point beyond which additional samples no longer improve data efficiency. By analogy, one can think of a critical evaluation budget $E_{\rm crit}$, and choose $E\approx E_{\rm crit}$, to optimally balance variance reduction against responsiveness. Our framework complements other recent monotonic improvement goals (Xie et al. 2025) without relying on stochastic density ratios, hence applies uniformly to both stochastic and deterministic settings.
@@ A.1   Concentration Inequality (Proof of Lemma 1)
Applying Hoeffding’s inequality to the bounded, independent variables $\{X_{i}\}$ (Hoeffding 1994) gives for any $\epsilon>0$ the one‑sided tail bounds
@@ A.4   Smoothness and Descent Properties
Follows from standard policy gradient smoothness results (e.g. Lemma D.3 in (Agarwal et al. 2021)), using the bounded Fisher information in Assumption 1. ∎
@@ A.5   Proof of Theorem 1
so $\min_{t<T}\|\nabla J(\pi_{\theta_{t}})\|=O(T^{-1/4})$. Lemma 41̇ of(Agarwal et al. 2021) then gives
@@ A.6   Proof of Theorem 2 (Biased Critic Case)
where $\epsilon_{t}$ is the bias that we assume is also upper bounded as well as its gradient by some $\epsilon\geq 0$, i.e., the estimation of the value function is accurate in both zeroth and first order. Such an assumption can be satisfied when a deep neural network is used to estimate value functions, e.g., (Du et al. 2019; Neyshabur 2017; Miyato et al. 2018), and is widely assumed in actor-critic analysis, e.g., (Wang et al. 2019a; Zhou and Lu 2023; Chen et al. 2023; Zhang et al. 2020; Qiu et al. 2021; Kumar et al. 2023; Xu et al. 2020b, a; Suttle et al. 2023).
@@
(Lemma D.3 of (Agarwal et al. 2021)) The value functions are $L$-smooth, i.e., there exists some constant $L$, such that
@@
Following a similar vanilla SGD analysis (Dutta et al. 2020, 2023), we obtain the results.
@@ B.4   Additional Experiments
In the main paper (Figure 2), we showed four continuous‐control suites for PPO, TRPO, and DDPG. Here, we complete the picture by adding results on two Box2D tasks (CarRacing, BipedalWalker) and two high‐DOF Bullet tasks (HalfCheetah, Humanoid). We also evaluated our profiling wrapper on the state-of-the-art off-policy actor-critic, TD3 (Fujimoto et al. 2018), that extends DDPG with clipped-double Q-learning, delayed policy updates, and target-policy smoothing. All of TD3’s core improvements—reduced overestimation bias from twin critics, more accurate actor gradients via delayed updates, and smoother targets via policy noise—are preserved inside our wrapper while reward profiling yields similar stability gains across both on‐ and off‐policy methods.